An amount of P340000 is invested, part at 8% and the remainder at 9%. Find the amount invested at each rate if the yearly income on each investment is the same.
Answer by Theo(13342)

x = part invested at 8%.
y = part invested at 9%

x + y = 340,000

.08x = .09y

solve for x to get x = .09y/.08 which becomes x = 1.125y

in the equation x + y = 340,000, replace x with 1.125y to get 1.125y + y = 340,000.

combine like terms to get 2.125y = 340,000.

solve for y to get y = 340,000 / 2.125 = 160,000.

since x = 1.125y, you get x = 1.125 * 160,000 = 180,000.

the amount invested at 8% is 180,000.

the amound invested at 9% is 160,000

.08 * 180,000 = 14,400.

.09 * 160,000 = 14,400.

the income is the same, so the solution is good.